OpenStudy (anonymous):

solve using substitution 1) 6x + 2y = 0 4x + y = -1 2) 3x + 2y = -4 x + 2y = -8 3) 5x - 4y = 14 x - 2y = -2 4) x + 3y = -4 4x + 5y = 5 5) x + 3y = -5 2x - y = 18 just answers please im running outa time!

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

i know i posted again so itall go back to the top! :P

OpenStudy (anonymous):

and umm...working out going to the gym and eating right y?

OpenStudy (anonymous):

i would just like help wwith my math please!

OpenStudy (jdoe0001):

lol

OpenStudy (anonymous):

yeah i thought this was a school/ homework help site not a creep on people site!

OpenStudy (jdoe0001):

let's do the 1st one, no that abb0t didn't already go through it 6x + 2y = 0 4x + y = -1 can you solve for "y" in the 2nd one?

OpenStudy (anonymous):

y is -1 right? abbOt did go over it but i just need answers now im running outa time and im quite frustrated with this!

OpenStudy (abb0t):

@love_jessika15

OpenStudy (jdoe0001):

well, \(\bf y \ne -1\)

OpenStudy (anonymous):

so 4x-1=-1?

OpenStudy (jdoe0001):

well, you're supposed to isolate "y" on say, the left-hand side

OpenStudy (anonymous):

so -1y=4x-1?

OpenStudy (jdoe0001):

4x + y = -1 # subtract 4x from both sides y = -1 -4x \(\bf 6x + 2y = 0\\ 4x + y = -1\\ \color{red}{y = -1 -4x}\\ \textit{let's use THAT "y" in the 1st equation}\\ 6x + 2y = 0 \implies 6x + 2(\color{red}{-1 -4x}) = 0\)

OpenStudy (jdoe0001):

so, from there, just solve for "x"